# Testing LRD in the spectral domain for functional time series in manifolds

### v3

@@ 4.1 Practical implementation
The practical implementation of the proposed statistical testing procedure, in terms of the Gaussian random projection methodology (see Theorem 4.1 in Cuesta-Albertos, Fraiman and Ransford (2007)), is now briefly discussed. The Karhunen–Loéve expansion in Lemma 4.6 below can be applied in such an implementation. Let us consider the random Fourier coefficients
where integration is understood in the mean–square sense, and is the limit Gaussian random element in introduced in Lemma 2.1.
Lemma 4.6.
Let be defined, as before, satisfying equation (23) in Lemma 2.1. Then, the following series expansion holds in the mean–square sense: For every
@@
Theorem 4.5 motivates the methodology to be adopted in practice. Specifically, as illustrated in the simulation study undertaken in the next section, a consistent test for LRD is obtained by rejecting when, for every
@@
The associated dimensionality problem can be substantially alleviated if we restrict our attention to the case where all moments of are finite and satisfy the Carleman condition under In that case, Theorem 4.1 in Cuesta-Albertos, Fraiman and Ransford (2007) leads to the following test statistic, evaluated conditionally to the observed functional value of a functional Gaussian random variable in the space whose probability measure on is denoted as Specifically, consider
Then, will be rejected if the observed value of is larger than an upper tail standard normal critical value. Note that, if holds then also holds, and if fails then also fails –a.s. Thus, with probability one, we will generate a realization of random direction in for which fails (see also Cuesta-Albertos et al. (2019)).
For one can consider the Gaussian random directions in
where denotes a zero–mean Gaussian random variable with variance Alternatively to the Gaussian random spherical harmonics considered in Marinucci, Rossi and Wigman (2020), for each we consider here the random kernel given by, for
Note that when and/or are dominant eigenspaces of the Laplace Beltrami operator on where the projected process displays the largest dependence range, the random kernel plays a key role in our approach. Furthermore, if we assume that the functional random variable can be generated from the Karhunen–Loéve expansion in Lemma 4.6. Thus, is a zero–mean Gaussian random element in satisfying (35), with covariance operator having kernel
for every
@@ 5 Simulation study
Our simulations will be set on An alternative generation algorithm to the ones considered in Ovalle–Muñoz and Ruiz–Medina (2024a) and Ovalle–Muñoz and Ruiz–Medina (2024b) is implemented reducing computational burden allowing for the consideration of large functional sample sizes. Theorem 2.2 is illustrated in the context of SPHARMA(p,q) processes, and the illustration of Theorem 4.5 is carried out in the context of multifractionally integrated SPHARMA(p,q) processes. These numerical results are respectively reported in Sections 5.1 and 5.2 for i.e., (see also Section 6 for a wider analysis for different values). When Section 5.3 opens new research lines beyond condition (16), providing empirical evidence of a faster a.s. divergence rate of in the norm of the space Finally, Section 5.4 shows empirical size and power properties of the testing procedure.
@@ 5.1 Asymptotic Gaussian distribution of under
Let us consider that the elements of the family of spectral density operators of have frequency varying eigenvalues, with respect to the system of eigenfunctions of the Laplace–Beltrami operator, obeying the following equation under (see Ruiz–Medina (2022)):
@@
Here, and are the so–called autoregressive and moving average operators, respectively. Also, for each and have not common roots, and their roots are outside of the unit circle (see also Corollary 6.17 in Beran (2017)). Thus, satisfies an SPHARMA(p,q) equation (see also Caponera (2021); Caponera and Marinucci (2021)).
@@
For each the empirical distribution of the centered and standardized projection into of is displayed in Figure 2 for functional sample size and repetitions, in Figure 3 for functional sample size and repetitions, and in Figure 4 for functional sample size and repetitions. These empirical distributions approximate the support and shape of a standard Gaussian probability density. The empirical standardization displays a decreasing pattern over the spherical scale meaning that the respective limit one–dimensional Gaussian probability measures of these projections have decreasing support. According to Theorem 1.2.1 in Da Prato and Zabczyk (2002), this property is satisfied by the infinite product Gaussian measure on whose restriction to is identified in the –sense with the probability measure of the limit Gaussian random element in Theorem 2.2.
@@ 5.2.1 Example 1
Theorem 4.5 is now illustrated in the case where the largest dependence range is displayed by the process projected into the eigenspace Figure 5 displays a sample realization of the corresponding multifractionally integrated SPHARMA (1,1) process projected into
@@ 5.3 Example 4
Under this scenario, conditions (16), and (32) in Theorem 4.2, are not satisfied. Indeed, we are out of the scenario where the summability in time of the square of the Hilbert–Schmidt operator norms of the elements of the covariance operator family holds. Then, new technical tools are required to address the asymptotic analysis in the spectral domain of this family of manifold supported functional time series displaying stronger levels of persistency in time. Let us again consider for projected into In this example, the multifractional integration of SPHARMA(1,1) process generated in Section 5.1 has been achieved in terms of LRD operator having eigenvalues displayed at the left–hand side of Figure 12, with and and being the dominant subspace. The same functional sample sizes as in Examples 1–3 have been considered. One can observe, in the three plots displayed at the right–hand side of Figure 12 , that the decreasing patterns over displayed in Figures 6, 8 and 10 do not hold in this example. Table 2 also illustrates a faster increasing than in Examples 1–3 of for functional sample sizes ranging from to under and
@@ 5.4 Empirical size and power analysis
6 Final comments and discussion
The simulation study illustrates four key aspects of our approach, briefly summarized in points (i)–(iv) below:
@@ 6 Final comments and discussion
allowing the application of Prokhorov Theorem to prove the convergence, as of to in the space In particular, the asymptotic Gaussian distribution of under follows. This result is illustrated in Section 5.1 from Theorem 1.2.1 in Da Prato and Zabczyk (2002).
The crucial role played by our design of the test statistic operator in the derivation of the conditions assumed to obtain consistency of the test (see Proposition 4.1 and Theorem 4.2). The simulation study also reveals that the additional conditions assumed in Theorem 4.5 lead to universal a.s. divergence rates, not affected by the localization of the dominant eigenspace, or the value of the parameter chosen, under the bandwidth parameter scenario when
@@
- (iv)
Regarding topic (ii), our final conclusions are also supported by the numerical results showed in Table 5, where, in the three examples analyzed, for large functional sample sizes the sample values of are displayed, when is projected into and
| T | Example 1 | Example 2 | Example 3 | ||
| 1.6885(1.0e+04) | 1.6384(1.0e+04) | 1.6438(1.0e+04) | |||
| 4.3549(1.0e+07) | 4.2251(1.0e+07) | 4.2290(1.0e+07) | |||
| 1.7493(1.0e+08) | 1.6944(1.0e+08) | 1.6989(1.0e+08) | |||
| 1.8067 (1.0e+04) | 1.7733(1.0e+04) | 1.7693(1.0e+04) | |||
| 4.4789(1.0e+07) | 4.3747(1.0e+07) | 4.3510(1.0e+07) | |||
| 1.7984(1.0e+08) | 1.7470(1.0e+08) | 1.7476(1.0e+08) | |||
| 2.0296(1.0e+04) | 2.0109(1.0e+04) | 1.9993(1.0e+04) | |||
| 4.5133(1.0e+07) | 4.4271(1.0e+07) | 4.4138(1.0e+07) | |||
| 1.8040 (1.0e+08) | 1.7518(1.0e+08) | 1.7624(1.0e+08) |
Although, as expected, the sample values of slightly increase when increases in Examples 1–3, no significant differences are observed in the sample divergence rate of for the three values of parameter analyzed in Table 5 (see the proof of Theorem 4.5). Furthermore, as proved in Theorem 4.5, under condition when one can also observe that the location of the dominant eigenspace does not affect the sample divergence rate. Note also that, as illustrated in Ovalle–Muñoz and Ruiz–Medina (2024a), the regular spectral factor corresponding here to the SPHARMA(p,q) functional spectrum, has not effect in the asymptotic analysis when its multifractionally integration is considered.
@@
Finally, we remark that all computations involved in the simulation study undertaken, in particular, in the implementation of the proposed inference tools and testing approach, have been achieved in terms of a unique orthonormal basis, given by the eigenfunctions of the Laplace Beltrami operator. Thus, we have worked under the scenario where the eigenfunctions of the elements of the covariance and spectral density operator families are known. This fact constitutes an important advantage of the analyzed setting, avoiding the use of empirical eigenfunction bases. We have also worked under the context of fully observed functional data. Under a manifold random uniform sampling design, the case of sparse discretely observed functional data can be addressed in a similar way, computing projections via numerical Monte–Carlo integration (see, e.g., Zhou, Wei and Yao (2024)).
@@
This work has been supported in part by projects PID2022–142900NB-I00 and PID2020-116587GB-I00, financed by MCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033), as well as supported by grant ED431C 2021/24 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).
@@ Appendix A Proofs of the results in Section 2
A.1 Proof of Theorem 2.2
Proof.
From Lemma 2.1,
@@
Under from Theorem D2 in the Supplementary Material of Panaretos and Tavakoli (2013a), keeping in mind (46),
@@
under SRD, and for certain and with as As before, denotes the weighted periodogram operator with kernel In equation (49), we have considered sufficiently large to apply the identity for and (see Lemma F11 in the Supplementary Material of Panaretos and Tavakoli (2013a)). Thus, under assuming the conditions in Lemma 2.1, the sequence is tight. Hence, the convergence, as of to the Gaussian random operator with kernel (see equation (45)), in the norm of the space follows from Prokhorov Theorem. Here, denotes the space of zero–mean second–order –valued random variables with the norm
Denote
@@ Appendix B Proofs of the results in Section 3
B.1 Proof of Lemma 3.1
Proof.
Let us consider
@@
From (57) and Parseval identity (see equation (59)),
From equations (60) and (61), Dominated Convergence Theorem then leads to
and Hence, the desired result follows from Parseval identity.
B.2 Proof of Corollary 3.2
@@
Applying Lemma 3.1, and Lemmas F10 and F12 of Appendix F in the Supplementary Material of Panaretos and Tavakoli (2013a),
@@
Under Assumption I, there exists an orthonormal basis of such that (see Helgason (1959))
@@
in the norm of the space where denotes the cumulant spectral density operator of order of under introduced in equation (30). ∎
@@ Appendix C Proofs of the results in Section 4
C.1 Proof of Proposition 4.1
Proof.
Under we have for every From Lemma 3.1, considering sufficiently large,
@@
where denotes the space of bounded linear operators on ∎
C.2 Proof of Theorem 4.2
Proof.
From Lemmas 3.1 and 3.3, applying trace formula, for sufficiently large,
@@
C.3 Proof of Corollary 4.4
Proof.
Applying triangle and Jensen inequalities, we obtain from Corollary 3.2 and Theorem 4.2,
C.4 Proof of Theorem 4.5
Proof.
The proof of this result shares some ideas with the proof of Theorem 2 of Harris et al. (2008), formulated in the time domain for real–valued time series. Specifically, the test statistic operator is rewritten as
@@
C.5 Proof of Lemma 4.6
Proof.
For consider a Gaussian random element in the space with kernel satisfying (23). Hence, from (47),

### v4

@@ 4.1 Practical implementation
The practical implementation of the proposed statistical testing procedure, in terms of the Gaussian random projection methodology (see Theorem 4.1 in [8]), is now briefly discussed. The Karhunen–Loéve expansion in Lemma 4 below can be applied in such an implementation. Let us consider the random Fourier coefficients
where integration is understood in the mean–square sense, and is the limit Gaussian random element in introduced in Lemma 1.
Lemma 4
Let be defined, as before, satisfying equation (23) in Lemma 1. Then, the following series expansion holds in the mean–square sense: For every
@@
Theorem 3 motivates the methodology to be adopted in practice. Specifically, as illustrated in the simulation study undertaken in the next section, a consistent test for LRD is obtained by rejecting when, for every
@@
The associated dimensionality problem can be substantially alleviated if we restrict our attention to the case where all moments of are finite and satisfy the Carleman condition. In that case, Theorem 4.1 in [8] leads to the following test statistic, evaluated conditionally to the observed functional value of a non–degenerated functional Gaussian random variable in the space whose probability measure on is denoted as Specifically, consider
Then, will be rejected if the observed value of is larger than an upper tail standard normal critical value. Note that, if holds then also holds, and if fails then also fails –a.s. Thus, with probability one, we will generate a realization of random direction in for which fails (see also [9]).
In the spirit of the Gaussian random degree– spherical harmonics introduced in [25] (), we will consider a truncated Karhunen–Loéve expansion in the generation of a non–degenerated Gaussian measure characterizing the random direction where our test statistics is projected (see equation (38)). Specifically, we consider a zero–mean Gaussian random element in with trace covariance operator having kernel
for every The trace property of can be equivalently expressed as Therefore, admits a series expansion, Karhunen–Loéve expansion (see, e.g., Lemma 4), whose truncated version is implemented. Particularly, one can construct centered isotropic Gaussian random fields by finite–dimensional projection
for and involving the Gaussian random directions in
where denotes a zero–mean Gaussian random variable with variance for Kernel plays a key role in our approach when and/or are dominant eigenspaces of the Laplace Beltrami operator on
@@ 5 Simulation study
Our simulations will be set on An alternative generation algorithm to the ones considered in [26] and [27] is implemented reducing computational burden allowing for the consideration of large functional sample sizes. Theorem 1 is illustrated in the context of SPHARMA(p,q) processes, and the illustration of Theorem 3 is carried out in the context of multifractionally integrated SPHARMA(p,q) processes. These numerical results are respectively reported in Sections 5.1 and 5.2 for i.e., (see also Section 6 for a wider analysis for different values). When Section 5.3 opens new research lines beyond condition (16), providing empirical evidence of a faster a.s. divergence rate of in the norm of the space Finally, Section 5.4 shows empirical size and power properties of the testing procedure.
@@ 5.1 Asymptotic Gaussian distribution of under
Let us consider that the elements of the family of spectral density operators of have frequency varying eigenvalues, with respect to the system of eigenfunctions of the Laplace–Beltrami operator, obeying the following equation under (see [36]):
@@
Here, and are the so–called autoregressive and moving average operators, respectively. Also, for each and have not common roots, and their roots are outside of the unit circle (see also Corollary 6.17 in [1]). Thus, satisfies an SPHARMA(p,q) equation (see also [5]; [6]).
@@
For each the empirical distribution of the centered and standardized projection into of is displayed in Figure 2 for functional sample size and repetitions, in Figure 3 for functional sample size and repetitions, and in Figure 4 for functional sample size and repetitions. These empirical distributions approximate the support and shape of a standard Gaussian probability density. The empirical standardization displays a decreasing pattern over the spherical scale meaning that the respective limit one–dimensional Gaussian probability measures of these projections have decreasing support. According to Theorem 1.2.1 in [10], this property is satisfied by the infinite product Gaussian measure on whose restriction to is identified in the –sense with the probability measure of the limit Gaussian random element in Theorem 1.
@@ 5.2.1 Example 1
Theorem 3 is now illustrated in the case where the largest dependence range is displayed by the process projected into the eigenspace Figure 5 displays a sample realization of the corresponding multifractionally integrated SPHARMA (1,1) process projected into
@@ 5.3 Example 4
Under this scenario, conditions (16), and (32) in Theorem 2, are not satisfied. Indeed, we are out of the scenario where the summability in time of the square of the Hilbert–Schmidt operator norms of the elements of the covariance operator family holds. Then, new technical tools are required to address the asymptotic analysis in the spectral domain of this family of manifold supported functional time series displaying stronger levels of persistency in time. Let us again consider for projected into In this example, the multifractional integration of SPHARMA(1,1) process generated in Section 5.1 has been achieved in terms of LRD operator having eigenvalues displayed at the left–hand side of Figure 12, with and and being the dominant subspace. The same functional sample sizes as in Examples 1–3 have been considered. One can observe, in the three plots displayed at the right–hand side of Figure 12 , that the decreasing patterns over displayed in Figures 6, 8 and 10 do not hold in this example. Table 2 also illustrates a faster increasing than in Examples 1–3 of for functional sample sizes ranging from to under and
@@ 5.4 Empirical size and power analysis
6 Final comments. Reliable inference from our approach
The simulation study illustrates six key aspects of our approach, briefly summarized in points (i)–(vi) below:
@@
allows the application of Prokhorov Theorem to prove the convergence, as of to in the space In particular, the asymptotic Gaussian distribution of under follows from this result. This result is illustrated in Section 5.1 from Theorem 1.2.1 in [10].
The crucial role played by the design of the test statistic operator in the derivation of the conditions assumed to obtain consistency of the test (see Proposition 1 and Theorem 2). The simulation study also reveals that the additional conditions assumed in Theorem 3 lead to universal a.s. divergence rates. These rates are not affected by the localization of the dominant eigenspace, or the value of the parameter chosen, under the bandwidth parameter scenario when Table 5 visualizes this fact for parameter values Specifically, in the three examples analyzed, the sample values of projected into are displayed under these three bandwidth parameter scenarios, for functional sample sizes Note that, although, as expected, the sample values of slightly increase when increases (see Examples 1–3), no significant differences are observed in the sample divergence rate of between the three values of parameter analyzed. Furthermore, under condition when one can observe the invariance of the sample divergence rate against the location of the dominant eigenspace. Keeping in mind that, as illustrated in [26], the regular spectral factor which here is represented by the SPHARMA(p,q) functional spectrum, has not impact in the asymptotic analysis, one can conclude the findings in Table 5 are representative.
Table 5: projected into T Example 1 Example 2 Example 3 1.6885(1.0e+04) 1.6384(1.0e+04) 1.6438(1.0e+04) 4.3549(1.0e+07) 4.2251(1.0e+07) 4.2290(1.0e+07) 1.7493(1.0e+08) 1.6944(1.0e+08) 1.6989(1.0e+08) 1.8067 (1.0e+04) 1.7733(1.0e+04) 1.7693(1.0e+04) 4.4789(1.0e+07) 4.3747(1.0e+07) 4.3510(1.0e+07) 1.7984(1.0e+08) 1.7470(1.0e+08) 1.7476(1.0e+08) 2.0296(1.0e+04) 2.0109(1.0e+04) 1.9993(1.0e+04) 4.5133(1.0e+07) 4.4271(1.0e+07) 4.4138(1.0e+07) 1.8040 (1.0e+08) 1.7518(1.0e+08) 1.7624(1.0e+08) -
@@
The testing approach adopted shows good empirical size and power properties for finite functional samples, as reported in Tables 3 and 4. Namely, Table 3, in the simulation study undertaken, shows empirical test sizes very close to the theoretical value for the minimum sample size considered, and for the number of repetitions Table 4 displays, for empirical powers in the interval at the eight random directions tested, and for the number of repetitions Summarizing, as illustrated in Section 5.4, for relatively small functional sample sizes, reliable inference based on our functional spectral nonparametric approach is possible. On the other hand, asymptotic properties like consistency of the test are verified to hold, as given in Table 1, displaying the increasing order of magnitude of the Hilbert–Schmidt operator norm of our test statistics, which is around for the smallest functional sample size in all examples. By the same reasons explained in (ii), i.e., the invariance of the results displayed in Tables 3 and 4 against the location of the dominant eigenspace ( and respectively in Examples 1,2,3), and the absence of asymptotic impact of the choice of the SPHARMA(p,q) functional spectrum, one can conclude the representativeness of the numerical results reflected in Tables 3 and 4.
@@
We remark that all computations involved in the simulation study undertaken, in particular, in the implementation of the proposed inference tools and testing approach, have been achieved in terms of a unique orthonormal basis, given by the eigenfunctions of the Laplace Beltrami operator. Thus, we have worked under the scenario where the eigenfunctions of the elements of the covariance and spectral density operator families are known. This fact constitutes an important advantage of the analyzed setting, avoiding the use of empirical eigenfunction bases. We have also worked under the context of fully observed functional data. The case of sparse discretely observed and contaminated functional data can be addressed from the nonparametric series least–squares estimation of our functional data set, and plug–in implementation of our test statistics (see, e.g., [46] and [37]). Specifically, under suitable restrictions on the local Hölder regularity of our functional data set, and on the supremum norm of the sieve basis elements, as well as on the pure point spectral properties of the autocovariance matrix of the random sieve basis (involved in the nonparametric series least–squares estimation of the discretely observed functional data set), one can derive similar asymptotic results. To this aim, a suitable manifold random uniform sampling design must be considered. The time–varying manifold sampling frequency must display a faster divergence rate than the time–varying sieve basis dimension, but slower than the functional sample size. Under these conditions, similar results on asymptotic bias analysis can be obtained under depending on the almost surely uniform convergence rates to the theoretical values of our functional data, involved in their nonparametric series least–squares estimation.
As commented in the Introduction, the presented approach allows the statistical inference from spatiotemporal data sets embedded into the sphere. Thus, the non–euclidean spatial statistical analysis of such data sets can be performed enhancing geometrical interpretability, avoiding the usual transformations of longitudes and latitudes of the data to work in a cartesian reference coordinate system. Under the invariance properties assumed in our setting, the case of discretely observed functional data can be addressed considering the sieve basis constructed from the eigenfunctions of the Laplace Beltrami operator. This sieve basis allows an easier geometrical implementation (see, e.g., [18], and [46] for alternative sieve bases in nonparametric series least–squares ridge regression in an euclidean spatial setting). An important dimension reduction is obtained in terms of this sieve basis, which is crucial in the reconstruction of high–dimensional data sets. This fact constitutes another remarkable feature of our approach that reduces computational burden, allowing the implementation of resampling techniques in real data applications. Note that temporal information can be incorporated under our functional time series framework, extending recent developments in the purely spatial statistical euclidean context (see, e.g., [18]). Just to mention a motivating data example for implementation of our approach, the authors in [18] analyze a georeferenced spatiotemporal discretely observed population data set to predict nighttime population in Tokyo. They implement series spatial ridge regression estimation after monthly averaging the data, ignoring time information that is crucial in this prediction problem. Our functional spectral nonparametric approach allows time information to be processed in an efficient way in an non–euclidean setting as commented in (iv) and (v).
@@
Acknowledgements. This work has been supported in part by projects PID2022–142900NB-I00 and PID2020-116587GB-I00, financed by MCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033), as well as supported by grant ED431C 2021/24 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).
@@ Appendix A Proofs of the results in Section 2
A.1 Proof of Theorem 1
Proof. From Lemma 2.1,
@@ A.1 Proof of Theorem 1
under SRD, and for certain and with as As before, denotes the weighted periodogram operator with kernel In equation (49), we have considered sufficiently large to apply the identity for and (see Lemma F11 in the Supplementary Material of [28]). Thus, under assuming the conditions in Lemma 2.1, the sequence is tight. Hence, the convergence, as of to the Gaussian random operator with kernel (see equation (45)), in the norm of the space follows from Prokhorov Theorem. Here, denotes the space of zero–mean second–order –valued random variables with the norm
Let us consider
@@ Appendix B Proofs of the results in Section 3
B.1 Proof of Lemma 2
Proof. Let us consider
@@ B.1 Proof of Lemma 2
B.2 Proof of Corollary 1
@@
Applying Lemma 3.1, and Lemmas F10 and F12 of Appendix F in the Supplementary Material of [28],
@@ B.3 Proof of Lemma 3
Under Assumption I, there exists an orthonormal basis of such that (see [17])
@@
in the norm of the space where denotes the cumulant spectral density operator of order of under introduced in equation (30).
@@ Appendix C Proofs of the results in Section 4
C.1 Proof of Proposition 1
Proof. Under we have for every From Lemma 3.1, considering sufficiently large,
@@ C.1 Proof of Proposition 1
where denotes the space of bounded linear operators on
C.2 Proof of Theorem 2
Proof. From Lemmas 3.1 and 3.3, applying trace formula, for sufficiently large,
@@ C.2 Proof of Theorem 2
C.3 Proof of Corollary 2
Proof. Applying triangle and Jensen inequalities, we obtain from Corollary 3.2 and Theorem 4.2,
C.4 Proof of Theorem 3
Proof. The proof of this result shares some ideas with the proof of Theorem 2 of [16], formulated in the time domain for real–valued time series. Specifically, the test statistic operator is rewritten as
@@ C.4 Proof of Theorem 3
C.5 Proof of Lemma 4
Proof. For consider a Gaussian random element in the space with kernel satisfying (23). Hence, from (47),